# Majority or Minority: Data Imbalance Learning Method for Named Entity Recognition

### v2

@@ 3.2 NER Frameworks and Loss Functions
Table 1: Performance with BERT in sequential labeling (in macro F1). In all items, MoM had the best score, with a significant difference from FL, which was the next best ($\alpha=0.05$).
CoNLL2003 OntoNotes5.0 KWDLC NER Wiki
@@
Table 2: Comparison of performance in each entity in sequential labeling of CoNLL2003.
Table 3: Summary of the performance of MRC on CoNLL2003.
@@
Prec. Rec. F1 Prec. Rec. F1
MISC 79.18 84.58 81.78 79.21 84.04 81.54
LOC 92.91 93.50 93.20 92.60 93.49 93.04
ORG 89.87 93.17 91.54 89.90 92.70 91.27
PER 97.66 97.88 97.77 97.76 97.55 97.65
$\mathcal{O}$ 99.72 99.28 99.50 99.69 99.31 99.50
@@
Table 3: Summary of the performance of MRC on CoNLL2003.
3.3 Implementation Details
Models.
@@ 3.3 Implementation Details
We utilized pre-trained models; the input length of these models was determined by the maximum number of tokens in a sentence ($M=128$), with padding tokens ([PAD]) used for filling the remaining space to maintain a consistent length. We fine-tuned in 10 epochs using the Adam optimizer (Kingma & Ba, 2014) for each task. Sequential labeling: We used pre-trained BERT (Devlin et al., 2019) as the baseline model. We set $D$ to 768, a learning rate of $2\times 10^{-5}$, and a batch size of 64. MRC: We used BERT-MRC (Li et al., 2020a) as the baseline model and set a learning rate of $3\times 10^{-5}$ and a batch size of 32.
Hyperparameters.
@@
The tree-structured Parzen estimator (TPE) (Bergstra et al., 2011), implemented in the Bayesian optimization library Optuna (Akiba et al., 2019), was used to maximize the F1 score of the validation data. For the sequential labeling experiments, the hyperparameters of WCE-2 ($\beta$) and FL ($\gamma$) were explored in the predetermined range of 1.0–10.0 and 0.0–10.0, respectively, considering their papers (Lin et al., 2017; Li et al., 2020b). For the MRC experiments, we set the hyperparameters of FL $\gamma=3.0$ and those of DL $\epsilon=1.0$ and $\delta=0.01$, which is based on the DL hyperparameters carefully tuned in Li et al. (2020b). The MoM hyperparameter ($\lambda$) was explored in the predetermined range of 0.0–1.0 in both frameworks. Thus, in the sequential labeling, we set the MoM hyperparameter $\lambda$ for the datasets CoNLL2003, OntoNotes5.0, KWDLC, and NER wiki to 0.175, 0.125, 0.357, and 0.212, respectively, while for MRC CoNLL2003 was set $\lambda$ to 0.446.
4 Results
@@
Table 1 presents a comparison of the performances of each method using BERT in sequential labeling. We confirmed that MoM learning consistently outperforms other methods across all four datasets. In all datasets, the performance using MoM learning was significant at $\alpha=0.05$ against the next best method (FL). The results using WCE-1 and WCE-2 demonstrated poor performance compared to the baseline; the reason is discussed in Sec. 5.
@@
Table 3 presents a comparison between the baseline with and without MoM learning for each entity in CoNLL2003. The prefixes of the entity classes B and I are merged to show the average performance of the respective classes, resulting in nine classes (e.g., B-PER, I-PER, B-LOC and $\mathcal{O}$) becoming five classes (e.g., PER, LOC, and $\mathcal{O}$). We confirm that MoM learning improves the performance of entity classes without compromising the performance of the $\mathcal{O}$-class.
@@
Table 3 presents the performance with the CoNLL2003 dataset using the MRC. The results confirm MoM learning also demonstrated the best performance and was significant at $\alpha=0.05$, against the next best method (DL).
5 Discussion
@@
The most important factor in NER is the score of the entity classes, rather than the overall score, including the $\mathcal{O}$-class, as the prediction performance of NER generally concerns the score including the $\mathcal{O}$-class. In practical situations in which entities are extracted and utilized, the performances of the entity classes hold greater significance. Although MoM learning appears to be a marginal improvement, we confirm that MoM learning improves the performance of minor entity classes without sacrificing the performance of the major $\mathcal{O}$-class, regardless of language.
@@
For WCE, we attempted two methods (i.e., WCE-1 and 2); however, a poorer performance than the baseline CE was observed, highlighting the challenges posed by multiclass NER, with its inherent long-tail distribution. As evidenced by various ML tasks, conventional weighting methods struggle with the delicate design of loss functions dependent on specific datasets and tasks (Valverde et al., 2017; Jadon, 2020). Thus, the experiments in Table 1 highlight the difficulty of applying WCE weighting to the sequential labeling of NER.
@@
MoM learning was effective at sequential labeling and MRC, especially the latter, where we observe the role of MoM learning in monitoring the number of entities. For example, in the sentence “Estadio Santiago Bernabéu opened in 1974.”, “Estadio”, “Santiago”, and “Bernabéu” are assigned to the classes B-LOC, I-LOC, I-LOC, and the other words are assigned to the $\mathcal{O}$-class. When considering a basic sentence, it is highly likely that the word (“opened”) following the last word (“Bernabéu”) of the entity belongs to be I-LOC or $\mathcal{O}$-class because sequences of different entity words are extremely rare, such as LOCATION after PERSON. Because MoM learning focuses more on $\mathcal{O}$-class words, the model can learn whether the final word belongs to the I-LOC or $\mathcal{O}$-class. In other words, the MoM can monitor how many entity words are consecutive, which is a factor in the improved performance of the words at the end of the entity.
6 Conclusion
@@
In this paper, we have proposed a novel learning method, MoM learning, to address NER tasks characterized by a data imbalance with a long-tail distribution consisting of a single class with many samples (the majority class) and multiple classes with a few samples (the minority classes). MoM learning is a simple and effective method that suppresses misclassifications of majority as minority classes by incorporating the loss of samples in which ground truth is the majority class into the loss of conventional ML models. Evaluation experiments using four datasets (two each in English and Japanese) showed that MoM learning outperforms existing and even state-of-the-art methods in addressing data imbalances regardless of language or framework.

### v3

@@ IV-C Loss Functions
TABLE III: The performance with BERT in the sequential labeling task (in macro F1). In all items, the proposed MoM had the best score, with a significant difference in FL, which was the next best ($\alpha=0.05$).
CoNLL2003 OntoNotes5.0 KWDLC Stockmark NER Wiki
@@
TABLE IV: The performance with RoBERTa in the sequential labeling task (in macro F1) . In all items, the proposed MoM had the best score, with a significant difference in FL, which was the next best ($\alpha=0.05$) except for OntoNote $\dagger$
CoNLL2003 OntoNotes5.0 KWDLC Stockmark NER Wiki
Prec. Rec. F1 Prec. Rec. F1 Prec. Rec. F1 Prec. Rec. F1
RoBERTa (baseline) 89.93 91.56 90.74 88.24 90.00 89.11 77.11 81.66 79.31 81.07 84.45 82.72
w/ WCE-1 85.86 91.31 88.50 59.55 82.59 69.20 46.85 77.38 58.31 39.51 49.58 43.98
(-2.24) (-19.91) (-21.00) (-38.74)
w/ WCE-2 89.88 91.94 90.90 86.61 90.49 88.51 68.86 77.23 72.80 75.72 81.19 78.36
(+0.16) (-0.60) (-6.51) (-4.36)
w/ FL 90.97 91.11 91.04 88.35 90.09 89.22 80.12 82.44 81.28 81.41 85.07 83.24
(+0.30) (+0.11) (+1.93) (+0.52)
w/ MoM 91.11 91.27 91.19 88.34 90.16 89.25$\dagger$ 81.10 82.60 81.85 81.85 85.23 83.54
(proposed) (+0.45) (+0.14) (+2.54) (+0.82)
TABLE V: Comparison of performance in each entity in sequential labeling of the CoNLL2003 dataset. The prefixes B and I are marged for display purposes.
@@
Prec. Rec. F1 Prec. Rec. F1 Diff
MISC 79.18 84.58 81.78 79.21 84.04 81.54 +0.24
LOC 92.91 93.50 93.20 92.60 93.49 93.04 +0.16
ORG 89.87 93.17 91.54 89.90 92.70 91.27 +0.27
PER 97.66 97.88 97.77 97.76 97.55 97.65 +0.12
$\mathcal{O}$ 99.72 99.28 99.50 99.69 99.31 99.50 0.00
TABLE VI: Summary of the performance of the MRC task on the CoNLL2003 dataset. MoM learning was also significant in MRC compared to FL and DL, which are state-of-the-art methods.
CoNLL2003
@@
TABLE VII: The performance comparison by word count on CoNLL2003 test data in sequential labeling. # words indicate the number of words in each data, and # data indicates the number of data with that word count. To evaluate the sentence-level, we show the average accuracy set for every 5 words for the bin, calculated based on matching the labels of all words.
sentence-level Acc.
# words # data $\rho_{\mathcal{O}}$ [%] BERT w/ FL w/ MoM
1-5 874 64.32 89.36 89.17 89.79
6-10 1,190 77.59 87.22 88.26 87.81
11-15 310 80.05 81.43 82.15 82.58
16-20 244 89.12 88.01 88.49 89.69
21-25 254 87.12 83.02 84.44 84.55
26-30 204 87.02 80.05 79.74 79.62
31-35 175 86.65 72.99 75.78 76.50
36- 202 75.63 85.98 86.12 86.56
IV-D Implementation Details
Models
@@ IV-D Implementation Details
We utilized pre-trained models, as shown in Table II, where the input length of these models was determined by the maximum number of tokens in a sentence ($M=128$) with padding tokens ([PAD]) used for filling the remaining space to maintain a consistent length. We fine-tuned each task in 10 epochs using the Adam optimizer [33].
During sequential labeling, we used pre-trained BERT [14] and RoBERTa [34] as baseline models. In the latter part of the $D$-dimensional special classification ([CLS]) token of these models, we attached a new head of $D\times(M\times\mathcal{Y})$ two-layered fully connected neural networks, and the head was trained to minimize the cross-entropy loss for each token. As such, these models output the predicted probabilities of each class corresponding to the words in the text. We set $D$ to 768, a learning rate of $2\times 10^{-5}$ and a batch size of 64, respectively.
In MRC, we used BERT-MRC [17] as the baseline model, where the latter part of the [CLS] token, we used the same architecture as sequential labeling with changed $\mathcal{Y}$ to $2$. As such, this model outputs a range corresponding to an entity for a sentence and $\mathcal{Y}$ binary classification. We set a learning rate of $3\times 10^{-5}$ and a batch size of 32, respectively.
Hyperparameters
@@
Tree-Structured Parzen Estimator (TPE) [29], implemented in the Bayesian optimization library Optuna [30], to maximize the F1 score of the validation data. For the sequential labeling experiments, the hyperparameters of WCE-2 ($\beta$) and FL ($\gamma$) were explored in the pre-determined range of 1.0-10.0 and 0.0-10.0, respectively, considering their papers [15, 16]. For the MRC experiments, we set the hyperparameters of FL $\gamma=3.0$, and those of DL $\epsilon=1.0$ and $\delta=0.01$, which is based on the DL hyperparameters carefully tuned in  [16]. The hyperparameters of MoM ($\lambda$) were explored in the pre-determined range of 0.0-1.0 in both frameworks. We show the results of tuning the MoM hyperparameter $\lambda$ in Table II.
V Results
(a) macro-F1
(b) sentence-level
(c) word-level
Figure 2: The result of evaluating the performance with different numbers of training data on the CoNLL2003. Since the sampling methods can not be adapted while preserving the distribution of labels, we use random undersampling to reduce the dataset to 3/4, 2/3, 1/2, 1/3, 1/10, 6/100, and 3/100. MoM learning showed the best results even under low resources, regardless of the number of training data.
@@ V Results
Tables III and IV present a comparison of the performances of each method using BERT and RoBERTa in sequential labeling. We confirmed that MoM learning consistently outperforms other tested methods across all four datasets, regardless of these models. In all results (4 datasets $\times$ 2 models), the performance using MoM learning was significant at $\alpha=0.05$ against the other best method (FL). In the OntoNotes5.0 results with RoBERTa (marked as $\dagger$), MoM learning was slightly less significant, but it was significant compared to the baseline CE. The results using WCE-1 and WCE-2 demonstrated poor performance compared to the baseline, and the reasons for this are discussed in Section VI.
@@
Table V compares the baseline with/without MoM learning for each entity in the CoNLL2003. The prefixes of the entity classes B and I are merged to show the average performance of the respective classes, resulting in 9 classes (e.g., B-PER, I-PER, B-LOC and $\mathcal{O}$) becoming 5 classes (e.g., PER, LOC, and $\mathcal{O}$). We confirm that MoM learning improves the performance of entity classes without compromising the performance of the $\mathcal{O}$-class.
@@
Table VI presents the performance with the CoNLL2003 dataset using the MRC. The results confirm MoM learning also demonstrated the best performance and was significant at $\alpha=0.05$, against the other best method DL.
Table VII shows CoNLL2003 test performance in sequential labeling by word count. For sentence-level evaluation, the score is based on accuracy matching all word labels in the sentence, rather than conventional macro-F1. To reduce data variability, the bin is set to 5 words, and average scores are calculated. Thus, MoM learning improves performance regardless of word count compared to baseline CE and FL.
Figure 2 presents the results of evaluating the performance with different numbers of training data on the CoNLL2003 regarding macro-F1, sentence-level accuracy, and word-level accuracy, respectively. Since the sampling methods can not be adapted while preserving the distribution of labels, we use random undersampling to reduce the dataset to 3/4, 2/3, 1/2, 1/3, 1/10, 6/100, and 3/100. In those evaluations, MoM learning showed the best results even under low resources, regardless of the number of training data.
VI Discussion
@@
The most important factor in NER is the score of the entity classes, rather than the overall score, including $\mathcal{O}$-class, as the prediction performance of NER generally concerns the score including the $\mathcal{O}$-class. In practical situations in which entities are extracted and utilized, the performances of the entity classes hold greater significance. Although MoM learning appears a marginal improvement, we confirm MoM learning improves the performance of minor entity classes without sacrificing the performance of the major $\mathcal{O}$-class, regardless of the language and models.
@@
For WCE, we attempted two methods (i.e., WCE-1 and WCE-2); however, a poor performance compared to the baseline CE was observed, highlighting the challenges posed by multiclass NER with its inherent long-tail distribution. As evidenced by various ML tasks, conventional weighting methods struggle with the delicate design of loss functions dependent on specific datasets and tasks [35, 36]. Thus, the experiments in Tables III and IV highlight the difficulty of applying WCE weighting to the sequential labeling of NER.
@@
MoM learning was effective at sequential labeling and MRC, especially the latter, where we observe the role of MoM learning in monitoring the number of entities. For example, in the sentence “Estadio Santiago Bernabéu opened in 1974.”, “Estadio”, “Santiago” and “Bernabéu” are assigned the classes B-LOC, I-LOC, I-LOC, and other words are assigned the $\mathcal{O}$-class, respectively. when considering a basic sentence, it is highly likely that the word (“opened”) following the last word (“Bernabéu”) of the entity belongs to I-LOC or $\mathcal{O}$-class because sequences of different entity words are extremely rare, such as LOCATION after PERSON. Because MoM learning focuses more on $\mathcal{O}$-class words, the model can learn whether the final word belongs to the I-LOC or $\mathcal{O}$-class. In other words, the MoM can monitor the number of consecutive entity words, which is a factor in the improved performance of the words at the end of the entity.
Compared to FL, MoM learning improved performance on each score even when data was sampled, as shown in Figure 2. In addition, MoM learning improved performance more as the data size decreased. This contribution is because FL needs to set a hyperparameter $\gamma$ that depends on the difficulty of classification. Since sampling methods produce variations in classification difficulty, setting the hyperparameter appropriately was difficult and could not have contributed to the performance. Therefore, MoM learning consistently contributed to the performance improvement because it did not depend on the difficulty of classification, and $\rho_{\mathcal{O}}$ was almost the same regardless of the word count. In addition, since MoM learning contributes to sentence-level accuracy regardless of data size, when we assume an application, this contribution is expected to reduce human load, such as rechecking by annotators.
MoM learning is successful in that it does not depend on the word count or data size, but there is still room for improvement. Since MoM learning focuses on misclassification between $\mathcal{O}$ classes and entity classes, more is needed for misclassification between entity classes. Further improving the prediction performance is possible by introducing a method specialized for entity classes.
VII Conclusion
@@
In this paper, we proposed a novel learning method, MoM learning, to tackle the NER task characterized by data imbalance with a long-tail distribution, consisting of a single class with many samples (the majority class) and multiple classes with few samples (the minority classes). MoM learning is a simple and effective method that reduces misclassifications of majority as minority classes by incorporating the loss of samples whose ground truth is the majority class into the loss of conventional ML models. Evaluation experiments using four datasets (English and Japanese) showed that MoM learning outperforms existing and even state-of-the-art methods in addressing data imbalances regardless of language, model, data size, or framework, whether sequential labeling or MRC.